OpenStudy (anonymous):

20x^1/4 all over 5x^-1/4

OpenStudy (anonymous):

x^1/4 over x^-1/4 = x^1/2 so the problem is \[\frac{ 20\sqrt{x} }{ 5 }\] = \[4\sqrt{x}\]